OpenStudy (anonymous):

beta gamma functions :

OpenStudy (anonymous):

\[\int\limits_{0}^{\infty}(\sqrt{x}~\cdot~e^{-x^2}dx)\int\limits_{0}^{\infty}\frac{ e^{-x^2} }{\sqrt{x}}dx\]

OpenStudy (anonymous):

multiplication of two integrals so how do i go about?

OpenStudy (anonymous):

since i know the substitution

OpenStudy (anonymous):

so should i solve the one of the integral first n then multiply ?

OpenStudy (anonymous):

let x^2=t

ganeshie8 (ganeshie8):

\[\int\limits_{0}^{\infty}(\sqrt{x}~\cdot~e^{-x^2}dx)\int\limits_{0}^{\infty}\frac{ e^{-x^2} }{\sqrt{x}}dx = \int\limits_{0}^{\infty}\int\limits_{0}^{\infty}\sqrt{\frac{x}{y}}~\cdot~e^{-(x^2+y^2)}dx\, dy\]

OpenStudy (anonymous):

'.' how did u do? u took x=y?

OpenStudy (anonymous):

please elaborate Sir

ganeshie8 (ganeshie8):

scratch that, you want to express it in terms of gamma function right ? then x^2 = t looks good to me!

OpenStudy (anonymous):

okay after that i substituted it into both integrals so now how do i solve further? as separate integrals or..?

OpenStudy (anonymous):

please answer :3

ganeshie8 (ganeshie8):

\(x^2 = t \\ \implies 2xdx = dt \\\implies dx = \dfrac{1}{2}t^{-1/2}dt\) \[\int\limits_{0}^{\infty}\frac{ e^{-x^2} }{\sqrt{x}}dx \\~\\=\int\limits_{0}^{\infty} \dfrac{e^{-t}}{t^{1/4}}\, \dfrac{1}{2}t^{-1/2}dt \\~\\= \dfrac{1}{2} \int\limits_0^{\infty}t^{-3/4}e^{-t}\,dt \\~\\ = \dfrac{1}{2} \int\limits_0^{\infty}t^{-1/4-1}e^{-t}\,dt \\~\\=\dfrac{1}{2}\Gamma(1/4) \]

ganeshie8 (ganeshie8):

As you can see, its just algebraic manipulation. You just need to try and put the given integral in the gamma integral form : https://upload.wikimedia.org/math/6/0/3/603515ff4df2cb431387e92bb6419c66.png

ganeshie8 (ganeshie8):

And yes, doing them separately seems to work..

OpenStudy (anonymous):

i got the same as you got

OpenStudy (anonymous):

wait i got t^{3/4}

ganeshie8 (ganeshie8):

looks i have got t^{-3/4} ?

OpenStudy (anonymous):

oops sorry yea its 1/4

ganeshie8 (ganeshie8):

How about the other integral ?

OpenStudy (anonymous):

please wait

OpenStudy (anonymous):

same t^{1/4}

ganeshie8 (ganeshie8):

I'm not following you, if psble could you take a screenshot of your work and attach ?

OpenStudy (anonymous):

wait i got the same as the above work done by u

ganeshie8 (ganeshie8):

Okay good, what about the first integral ?

OpenStudy (anonymous):

so its getting squared

ganeshie8 (ganeshie8):

\[\int\limits_{0}^{\infty}(\sqrt{x}~\cdot~e^{-x^2}dx) = ?\]

OpenStudy (anonymous):

both turn out to be the same

OpenStudy (anonymous):

wait ill write it down

ganeshie8 (ganeshie8):

I'm getting \[\int\limits_{0}^{\infty}(\sqrt{x}~\cdot~e^{-x^2}dx) = \dfrac{1}{2}\Gamma(3/4)\]

ganeshie8 (ganeshie8):

Just double check your work..

OpenStudy (anonymous):

oops lol ROFL i was looking at the same integral!!!

OpenStudy (anonymous):

yup yup it is 3/4 so shall i leave it like that?

OpenStudy (anonymous):

@ganeshie8 ans=\(\huge\frac{\sqrt2\cdot \pi}{4}\)

ganeshie8 (ganeshie8):

Oh that looks nice, lets try and simplify maybe

OpenStudy (anonymous):

i believe this will also work usually we keep it till this ans

ganeshie8 (ganeshie8):

\[\int\limits_{0}^{\infty}(\sqrt{x}~\cdot~e^{-x^2}dx)\int\limits_{0}^{\infty}\frac{ e^{-x^2} }{\sqrt{x}}dx = \dfrac{1}{2}\Gamma(3/4)*\dfrac{1}{2}\Gamma(1/4)\]

OpenStudy (anonymous):

thats the final ans i got

ganeshie8 (ganeshie8):

Familiar with below relation between beta and gamma functions ? \[\beta(a,b)=\dfrac{\Gamma(a)\Gamma(b)}{\Gamma(a+b)}\]

ganeshie8 (ganeshie8):

Rearranging above identity a bit we get \[\Gamma(a)\Gamma(b) = \Gamma(a+b)*\beta(a,b)\]

ganeshie8 (ganeshie8):

try using that and see if it simplifies any further..

OpenStudy (anonymous):

i used another property :)

ganeshie8 (ganeshie8):

Yeah I have the feeling that there must be some smart way to work this... What property have you used ?

OpenStudy (anonymous):

\(\Gamma(n) \Gamma(1-n)=\Large \frac{\pi}{sin(n\pi)} \\ where 0<n<1\) oops the ans it should be sin(3\(\pi\)/4)